Deployment team & sites

EVO - GridPP Deployment team & sites meeting

EVO - GridPP Deployment team & sites meeting

Jeremy Coles
- This is the biweekly DTEAM & sites meeting - The intention is to run the meeting in EVO: Join the meeting in the GridPP Community area. - The phone bridge number is +44 (0)161 306 6802 (CERN number +41 22 76 71400). The phone bridge ID is 64688 with code: 4880.
    • 11:00 11:20
      Experiment problems/issues 20m
      Review of weekly issues by experiment/VO - LHCb -- Check on transfer problems - Only Brunel now seeing problems? - CMS - ATLAS - Other - Experiment blacklisted sites - Experiment known events affecting job slot requirements - Site performance issues -- September availability figures will be distributed shortly. We are aware of an issue whereby a problematic WMS has affected results.
    • 11:20 11:40
      Meetings & updates 20m
      - ROD team status (any points to raise to sites or issues to follow up?) - Tier-1 update - NGI support process... some recent problems in tickets submitted via GGUS reaching NGS/NGI colleagues who now support some of our services. This is being addressed but please raise any concerns quickly. One general comment, please provide as much information in the ticket as possible for the TPM to make a more accurate assessment. - The most recent EGI Grid Ops meeting was yesterday: Main item picked out by SP was: "Current deadline of 31.01.2011 for BDII on 3.1 means that the lcg-CE will loose support of critical component (resource BDII) then. As lcg-ce is still needed; so that's a problem". - The next WLCG Grid Deployment Board will be on 13th October (the agenda is here Graeme is the T2 rep this month but any of you can join via EVO (WLCG community). Are there any issues for us to raise? Topics to be covered are: Operational security - the recent vulnerability and response OPN troubleshooting Middleware - update on 3.1 decisions Installed capacity - starting to assume that is now accurate for capacity (not jobs). Then a series of talks on the Data Access & Management Demonstrators. - Escalated tickets 3 tickets on hold. 56316 - NGS-RAL not in BDII 58733 - RAL-PP biomed. dcache issue. 61228 - ALICE RAL T1.
    • 11:40 12:00
      Deployment priorities & timelines 20m
      - - CREAM Is this summary correct: Sites that have CREAM working are RAL T1, Glasgow, Manchester, Oxford, Birmingham and Imperial. Liverpool has CREAM setup but not in production. Installation is scheduled at RALPP, QMUL, UCL and Brunel. Further information is required for Lancaster, Sheffield, Cambridge, Bristol and Durham. - APEL - ARGUS
    • 12:00 12:02
      AOB 2m
      - New CA release - Next HEPSYSMAN meeting Monday 22nd November in Birmingham. - Top-level BDII decision in WLCG
    • 12:02 12:10
      Q32010 Tier-2 reports 8m
      - Observations & comments - Any escalation issues?